cn someone please help me find the x intercept?
y=x^2+x-6

Answer by doctor_who(15)

The "intercept" is where the graph crosses the x axis. In other words, the values for x which give you the result of y=0.
we have y = x^2+x-6
and we are looking for
0 = x^2+x+6 (since this will tell us where the intercepts are)
Now we just have to solve this equation.
A fast way to do this is to see that you can factor the right hand side of the equation like this :
0 = (x+3)(x-2)
Now in order for the expression to equal zero EITHER the stuff in the left hand set of brackets must equal zero OR the stuff in the right hand set of brackets must equal zero (since anything times zero equals zero)
So for this to be the case we have TWO solutions for the x intercept,
namely x = -3 (negative 3) or x = 2 (ANSWER)
Hope this helps !